OpenStudy (anonymous):

find the value of tan 5π/8

OpenStudy (amistre64):

well; we can either try to se trig formulas; or simply use a calculator

OpenStudy (anonymous):

oh, sorry. no calculators, says to use trig functions.

OpenStudy (amistre64):

you wanna use either a tan(a+b) or a tan(a-b) formula than; and find 2 angles that either add to or subtract to 5pi/8

OpenStudy (anonymous):

nice website, thanks

OpenStudy (amistre64):

tana+tanb tan(a+b) = ------------- 1- tana.tanb tana-tanb tan(a-b) = ------------- 1+ tana.tanb

OpenStudy (amistre64):

tan(2a) = 2.tana --------- 1-tan^2(a)

OpenStudy (amistre64):

but i dont see that one being helpful :)

OpenStudy (amistre64):

i dont se any of the familiar angles summing it up to get 112.5

OpenStudy (amistre64):

it is 225 degrees/2 tho; so maybe split that tan into sin and cos functions and work on half angle formulas

OpenStudy (amistre64):

either way you go, its gonna be a nightmare

OpenStudy (anonymous):

i know, but it looks like it would be so easy!

OpenStudy (anonymous):

thanks for trying

OpenStudy (amistre64):

it might be; but youd have to do allllooooottt of typing; and its just not that fun to try to keep track of on a keyboard

OpenStudy (amistre64):

sqrt(1+cos(5pi/4)) ----------------- = tan(5pi/8) sqrt(1-cos(5pi/4))

OpenStudy (amistre64):

sqrt(1+[-1/sqrt(2)]) ------------------- sqrt(1 - [1/sqrt(2)])

OpenStudy (amistre64):

\[\sqrt{\frac{\sqrt{2}-1}{\sqrt{2}+1}}\]

OpenStudy (amistre64):

thats the exact form

OpenStudy (amistre64):

well; i think it finally settles down to: -1 - sqrt(2)